OpenStudy (anonymous):

Differential Equation y'=3y^(2/3) and y(2)=0

OpenStudy (anonymous):

\[y'=3\sqrt[3]{y^2}\]

OpenStudy (anonymous):

\[y(2)=0\]

OpenStudy (anonymous):

the final answer is \[y=(x-2)^3\]

OpenStudy (anonymous):

if you guys could help me out with it it will literally

OpenStudy (anonymous):

make my day*

OpenStudy (tkhunny):

Have you considered the antiderivative? \(y = 3 \dfrac{y^{5/3}}{5/3} + C\) Simplify first, of course.

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

is it the same? can you explain?

OpenStudy (anonymous):

@shereenkhan ?

OpenStudy (anonymous):

nah @tkhunny I don't think that's the way tbh

rvc (rvc):

what do you have to do?

rvc (rvc):

to get y need to integrate y'

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

wby not just writing instead of y'=dy/dx?

rvc (rvc):